a women travels one mile due north, then two miles due east, then three miles due north again, and then once more east for 4 miles. how far is she from her starting point? (it is less than 5+ square root 5 miles.
Answer by checkley77(12844)

She is 4 miles north & 6 miles east of her starting location.
Thus you have a right triangle with sides of 4 & 6 miles and need to find the hypotenuse.
a^2+b^2=c^2
4^2+6^2=c^2
16+36=c^2
52=c^2
c=sqrt52
c=7.211 miles is her distance from her starting point.